OpenStudy (anonymous):

2/5x − 3 = 13 solve. By the way, the / stands for the fraction bar

OpenStudy (anonymous):

OK, just so we're clear, is that 2 over 5x or two x over 5?

OpenStudy (anonymous):

2/5x = 16 2 = 80x x = 2/80 = 1/40

OpenStudy (anonymous):

2 over 5 times x

OpenStudy (anonymous):

I can still take that either way \[\frac{2x}{5}\frac{2}{5x}\] First or second?